OpenStudy (anonymous):

Is there anyone who's bored and wants to try finding the inverse of a matrix?

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 2 \\ 5 & 9\end{matrix}\right]\]

OpenStudy (maheshmeghwal9):

\[\left[\begin{matrix}9 & -2 \\ -5 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

Is there a special process for two-by-two matrices?

OpenStudy (maheshmeghwal9):

yeah of course

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1 & 2 \\ 5 & 9\end{matrix}\right|\left.\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]=\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right|\left.\left(\begin{matrix}1 & 2 \\ 5 & 9\end{matrix}\right)^{-1}\right]\]

OpenStudy (maheshmeghwal9):

If \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]is a matrix then its inverse is : -\[\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (maheshmeghwal9):

gt it?

OpenStudy (unklerhaukus):

or use row operations to convert my LHS to my RHS

OpenStudy (anonymous):

Oh. I never knew that. Ok. It always took me forever for inverses of matrices. Thanks :) And I do know how to find the inverse using the normal way.

OpenStudy (maheshmeghwal9):

this is a special case for only order 2 matrix which i told u:)